OpenStudy (anonymous):

Find the radius of a circle with the equation x2 + y2 - 6x - 12y + 36 = 0

OpenStudy (welshfella):

you can do this by completing the square and tracform the equation to (x -a)^2 + (y - b)^2 = r^2

OpenStudy (welshfella):

* transform

OpenStudy (anonymous):

I'm not really sure how to do completing the square on this problem

OpenStudy (welshfella):

complete the square on x^2 - 6x;- Divide -6 by 2:- = (x - 3)^2 - 9

OpenStudy (welshfella):

and for y^2 - 12 y its (y - 6)^2 - 36

OpenStudy (welshfella):

x^2 + y^2 - 6x - 12y + 36 = 0 (x - 3^)2 + (y - 6( ^2 - 9 - 36 + 36 = 0 (x - 3)^2 + (y - 6)^2 = 9 compare this with the standard equation we have r^2 = 9 so the radius r = 3